6x-1/((x-3)(x+1)) into partial fractions
1
Expert's answer
2019-03-01T10:24:51-0500

"\frac{6x-1}{(x-3)(x+1)}=\frac{A}{x-3}+\frac{B}{x+1}=\frac{A(x+1)+B(x-3)}{(x-3)(x+1)}=\frac{Ax+A+Bx-3B}{(x-3)(x+1)}=\frac{x(A+B)+A-3B}{(x-3)(x+1)}"

"\begin{cases}
 A+B=6 \\
A-3B=-1
\end{cases}"

"\begin{cases}
 A+B=6 \\
A=3B-1
\end{cases}"

"3B-1+B=6"

"4B=7"

"B=\frac{7}{4}"

"A=3\cdot\frac{7}{4}-1=\frac{17}{4}"

"\frac{6x-1}{(x-3)(x+1)}=\frac{17}{4(x-3)}+\frac{7}{4(x+1)}"